# Enhancing Link Prediction with Fuzzy Graph Attention Networks and Dynamic Negative Sampling

### v1

@@ Abstract
I Introduction
@@ II Related Work
Graph Neural Networks have demonstrated remarkable versatility across various graph-based learning tasks. In node classification, seminal works like GraphSAGE [1] and Graph Attention Networks (GAT) [2] have established foundational approaches for learning node representations through neighborhood aggregation. GCN [3] introduced convolutional operations on graphs, enabling efficient feature propagation across network structures. For graph classification tasks, hierarchical pooling mechanisms have been developed, with DiffPool [4] and TopKPool [5] proposing learnable strategies to generate graph-level representations. In the context of link prediction, SEAL [6] pioneered the use of local subgraphs for edge existence prediction, while VGAE [7] employed variational autoencoders for learning edge formation patterns. Recent advances include NGNN [8], which introduces neural architecture improvements specifically designed for link prediction tasks.
@@
Fuzzy rough sets theory, initially proposed by Dubois and Prade [9], has evolved into a powerful framework for handling uncertainty and imprecision in data analysis. In feature selection, Jensen and Shen [10] developed fuzzy-rough attribute reduction algorithms that significantly outperform traditional approaches in identifying relevant features while maintaining information fidelity. The application of fuzzy rough sets in medical diagnosis has been exemplified by works such as [11], where they effectively handle the inherent uncertainty in patient data for more accurate disease classification. For uncertainty measurement, the framework has been extensively studied in theoretical works by [12] and [13], establishing mathematical foundations for quantifying various types of uncertainty in data relationships. Recent developments include hybrid approaches combining fuzzy rough sets with deep learning and applications in big data analytics [14] , demonstrating the framework’s adaptability to modern computational challenges.
@@
Except the innovative applications of Graph Neural Networks and fuzzy rough set theory in handling complex data and multi-task learning, against the backdrop of rapid advancements in machine learning, applications like knowledge graph technology in intelligent question-answering systems have become a key area of development for graph learning methods, effectively integrating multiple data sources to support flexible knowledge processing[15]. Similarly, multi-model integration technology applied to automated generation systems has significantly enhanced content generation flexibility and quality, providing valuable insights for representation learning based on graph data[16]. In network security, Graph Neural Networks have demonstrated strong generalization capabilities in supporting Botnet detection through machine learning, precisely identifying abnormal behaviors and enhancing network defense levels[17].
@@
In recent years, the scope of machine learning applications has continuously expanded, especially in image processing and VR fields. Blockchain-enhanced image retrieval systems, for example, have brought revolutionary improvements in data security and retrieval efficiency[18]. In VR and robotic interaction, AI-vision-powered intelligent systems have explored new methods for balancing real-world interaction and virtual immersion, making human-computer interaction more natural and seamless[19]. Furthermore, the integration of fuzzy rough set theory with deep learning has also achieved breakthroughs in traffic data prediction, enabling more accurate short-term forecasting through multi-source data fusion, effectively adapting to the dynamic demands of complex data environments[20]. Overall, these technological innovations not only showcase the broad applicability of graph learning methods and fuzzy rough sets across different task scenarios but also reinforce their theoretical and practical value in big data and intelligent applications.
III Methodology
@@
Figure 1: The FGAT Framework
III-A Fuzzy Negative Sampling
A fuzzy information system is defined as a tuple $(U,A,V,f)$, where $U$ represents a non-empty finite set of samples, $A$ denotes the finite set of sample attributes, $V$ represents the domain of all attributes in $A$, expressed as $V=\bigcup_{i}V_{i}$ where $V_{i}$ is the domain of attribute $i$, and $f$ is a mapping function $U\times A\rightarrow V$ [11].
For an attribute set $B\subseteq A$ and a fuzzy equivalence relation $R$, we can compute a coverage of the universe $U$. For a sample $x$, we denote its coverage under the fuzzy equivalence relation $R$ as $[x]_{R}$. The membership of a sample $y$ to the coverage $[x]_{R}$ is defined as $[x]_{R}(y)=R(x,y)$, where $R(x,y)$ quantifies the similarity between samples $x$ and $y$ under relation $R$. For any sample $x\in U$ and subset $X\subseteq A$, the fuzzy lower and upper approximations of sample $x$ to $X$ are defined as [11]:
$\begin{split}\underline{R_{S}}X(x)&=\inf_{y\in U}S(N(R(x,y)),X(y)),\\ \overline{R_{T}}X(x)&=\sup_{y\in U}T(R(x,y),X(y))\end{split}$ (1)
@@ III-A Fuzzy Negative Sampling
$\begin{split}\underline{R_{B}}d_{i}(x)&=\inf_{y\in U}\max(1-R(x,y),d_{i}(y)),\\ \overline{R_{B}}d_{i}(x)&=\sup_{y\in U}\min(R(x,y),d_{i}(y))\end{split}$ (2)
@@
$Score(x,y)=\alpha\times\underline{R_{B}}d_{y}(x)+(1-\alpha)\times\underline{R_{B}}d_{x}(y)$ (3)
@@
III-B FGAT Convolution Layer
@@
$e_{vu}=\text{LeakyReLU}\left(\mathbf{a}^{T}\left[\mathbf{W}\mathbf{h}_{v}\parallel\mathbf{W}\mathbf{h}_{u}\right]\right)$ (4)
@@
$\alpha_{vu}=\frac{\exp(e_{vu})}{\sum_{k\in\mathcal{N}(v)}\exp(e_{vk})}$ (5)
@@
$\mathbf{h}_{v}^{\prime}=\sigma\left(\sum_{u\in\mathcal{N}(v)}\alpha_{vu}\mathbf{W}\mathbf{h}_{u}\right)$ (6)
@@
$\mathbf{h}_{v}^{\prime}=\parallel_{k=1}^{K}\sigma\left(\sum_{u\in\mathcal{N}(v)}\alpha_{vu}^{(k)}\mathbf{W}^{(k)}\mathbf{h}_{u}\right)$ (7)
@@
Layer normalization [21] is incorporated to stabilize and expedite the training process by normalizing layer inputs. For an input vector $\mathbf{h}=[h_{1},h_{2},\dots,h_{d}]$ with $d$ features, the normalized output $\mathbf{\hat{h}}=[\hat{h}_{1},\hat{h}_{2},\dots,\hat{h}_{d}]$ is computed as:
$\hat{h}_{i}=\frac{h_{i}-\mu}{\sqrt{\sigma^{2}+\epsilon}}$ (8)
@@
$y_{i}=\gamma\hat{h}_{i}+\beta$ (9)
III-C The FGAT Framework
@@
$P_{r}^{link}(x,y)=\text{Sigmoid}(h_{x}h_{y}^{T})$ (10)
@@
IV Experiments
@@
IV-A Datasets
Our evaluation utilizes two research collaboration network datasets, summarized in Table I.
TABLE I: Datasets Summary
@@ IV-A Datasets
TABLE II: Experiment Results
Methods Ca-netscience Ca-sandi-auths
Precision Recall F1 ROC Precision Recall F1 ROC
MLP 0.4931 0.5879 0.5363 0.5044 0.5581 1.0000 0.7164 0.6181
GCN 0.5903 0.7363 0.6553 0.7078 0.5417 0.5417 0.5417 0.5304
GraphSAGE 0.5502 0.8132 0.6563 0.6277 0.4651 0.8333 0.5970 0.5885
GAT 0.6034 0.7692 0.6763 0.6916 0.6053 0.9583 0.7419 0.7240
FGAT 0.6667 0.6593 0.6630 0.7422 0.6216 0.9583 0.7541 0.7170
@@
IV-B Experiment Settings and Evaluation Metrics
We benchmark FGAT against several prominent baseline models: MLP, GCN [3], GraphSAGE [1], and GAT [2]. All baseline models maintain their default parameter configurations. For FGAT implementation, we configure the embedding dimension to 128 and employ a stack of 4 FGAT convolution layers. The dataset partitioning follows the aforementioned 0.7:0.1:0.2 ratio for training, validation, and testing, respectively.
@@ IV-B Experiment Settings and Evaluation Metrics
IV-C Results
The experimental results are presented in Table II, yielding several significant observations:

### v3

@@ Abstract.
conference: April 26–28, 2025; Tokyo, Japan; ccs: Computing methodologies Machine learningccs: Information systems Data miningccs: Information systems Information retrievalccs: Theory of computation Dynamic graph algorithms
1. Introduction
@@ 2. Related Work
Graph Neural Networks have demonstrated remarkable versatility across various graph-based learning tasks. In node classification, seminal works like GraphSAGE (hamilton2017inductive, ) and Graph Attention Networks (GAT) (velivckovic2017graph, ) have established foundational approaches for learning node representations through neighborhood aggregation. GCN (kipf2016semi, ) introduced convolutional operations on graphs, enabling efficient feature propagation across network structures. For graph classification tasks, hierarchical pooling mechanisms have been developed, with DiffPool (ying2018hierarchical, ) and TopKPool (diehl2019edge, ) proposing learnable strategies to generate graph-level representations. In the context of link prediction, SEAL (zhang2018link, ) pioneered the use of local subgraphs for edge existence prediction, while VGAE (kipf2016variational, ) employed variational autoencoders for learning edge formation patterns. Recent advances include NGNN (zhang2021nested, ), which introduces neural architecture improvements specifically designed for link prediction tasks.
@@
Fuzzy rough sets theory, initially proposed by Dubois and Prade (dubois1990rough, ), has evolved into a powerful framework for handling uncertainty and imprecision in data analysis. In feature selection, Jensen and Shen (jensen2004fuzzy, ) developed fuzzy-rough attribute reduction algorithms that significantly outperform traditional approaches in identifying relevant features while maintaining information fidelity. The application of fuzzy rough sets in medical diagnosis has been exemplified by works such as (xing2022weighted, ), where they effectively handle the inherent uncertainty in patient data for more accurate disease classification. For uncertainty measurement, the framework has been extensively studied in theoretical works by (gao2022parameterized, ) and (ye2021novel, ), establishing mathematical foundations for quantifying various types of uncertainty in data relationships. Recent developments include hybrid approaches combining fuzzy rough sets with deep learning and applications in big data analytics (ji2021fuzzy, ) , demonstrating the framework’s adaptability to modern computational challenges.
@@
Except the innovative applications of Graph Neural Networks and fuzzy rough set theory in handling complex data and multi-task learning, against the backdrop of rapid advancements in machine learning, applications like knowledge graph technology in intelligent question-answering systems have become a key area of development for graph learning methods, effectively integrating multiple data sources to support flexible knowledge processing (lin2021research, ). Similarly, multi-model integration technology applied to automated generation systems has significantly enhanced content generation flexibility and quality, providing valuable insights for representation learning based on graph data (yang2021research, ). In network security, Graph Neural Networks have demonstrated strong generalization capabilities in supporting Botnet detection through machine learning, precisely identifying abnormal behaviors and enhancing network defense levels (yang2022botnet, ).
@@
In recent years, the scope of machine learning applications has continuously expanded (cheng24patch, ; cheng25unifying, ; cheng22deep, ; cheng22estimation, ; xing24traffic, ; cheng23shapnn, ), especially in image processing and VR fields. Blockchain-enhanced image retrieval systems, for example, have brought revolutionary improvements in data security and retrieval efficiency (zhao2022image, ). In VR and robotic interaction, AI-vision-powered intelligent systems have explored new methods for balancing real-world interaction and virtual immersion, making human-computer interaction more natural and seamless (yang2023balancing, ). Furthermore, the integration of fuzzy rough set theory with deep learning has also achieved breakthroughs in traffic data prediction, enabling more accurate short-term forecasting through multi-source data fusion, effectively adapting to the dynamic demands of complex data environments (deng2021short, ). Overall, these technological innovations not only showcase the broad applicability of graph learning methods and fuzzy rough sets across different task scenarios but also reinforce their theoretical and practical value in big data and intelligent applications.
3. Methodology
@@
Figure 1. The FGAT Framework
3.1. Fuzzy Negative Sampling
A fuzzy information system is defined as a tuple $(U,A,V,f)$, where $U$ represents a non-empty finite set of samples, $A$ denotes the finite set of sample attributes, $V$ represents the domain of all attributes in $A$, expressed as $V=\bigcup_{i}V_{i}$ where $V_{i}$ is the domain of attribute $i$, and $f$ is a mapping function $U\times A\rightarrow V$ (xing2022weighted, ).
For an attribute set $B\subseteq A$ and a fuzzy equivalence relation $R$, we can compute a coverage of the universe $U$. For a sample $x$, we denote its coverage under the fuzzy equivalence relation $R$ as $[x]_{R}$. The membership of a sample $y$ to the coverage $[x]_{R}$ is defined as $[x]_{R}(y)=R(x,y)$, where $R(x,y)$ quantifies the similarity between samples $x$ and $y$ under relation $R$. For any sample $x\in U$ and subset $X\subseteq A$, the fuzzy lower and upper approximations of sample $x$ to $X$ are defined as (xing2022weighted, ):
(1) $\begin{split}\underline{R_{S}}X(x)&=\inf_{y\in U}S(N(R(x,y)),X(y)),\\ \overline{R_{T}}X(x)&=\sup_{y\in U}T(R(x,y),X(y))\end{split}$
@@ 3.1. Fuzzy Negative Sampling
(2) $\begin{split}\underline{R_{B}}d_{i}(x)&=\inf_{y\in U}\max(1-R(x,y),d_{i}(y)),\\ \overline{R_{B}}d_{i}(x)&=\sup_{y\in U}\min(R(x,y),d_{i}(y))\end{split}$
@@
(3) $Score(x,y)=\alpha\times\underline{R_{B}}d_{y}(x)+(1-\alpha)\times\underline{R_{B}}d_{x}(y)$
@@
3.2. FGAT Convolution Layer
@@
(4) $e_{vu}=\text{LeakyReLU}\left(\mathbf{a}^{T}\left[\mathbf{W}\mathbf{h}_{v}\parallel\mathbf{W}\mathbf{h}_{u}\right]\right)$
@@
(5) $\alpha_{vu}=\frac{\exp(e_{vu})}{\sum_{k\in\mathcal{N}(v)}\exp(e_{vk})}$
@@
(6) $\mathbf{h}_{v}^{\prime}=\sigma\left(\sum_{u\in\mathcal{N}(v)}\alpha_{vu}\mathbf{W}\mathbf{h}_{u}\right)$
@@
(7) $\mathbf{h}_{v}^{\prime}=\parallel_{k=1}^{K}\sigma\left(\sum_{u\in\mathcal{N}(v)}\alpha_{vu}^{(k)}\mathbf{W}^{(k)}\mathbf{h}_{u}\right)$
@@
Layer normalization (xiong2020layer, ) is incorporated to stabilize and expedite the training process by normalizing layer inputs. For an input vector $\mathbf{h}=[h_{1},h_{2},\dots,h_{d}]$ with $d$ features, the normalized output $\mathbf{\hat{h}}=[\hat{h}_{1},\hat{h}_{2},\dots,\hat{h}_{d}]$ is computed as:
(8) $\hat{h}_{i}=\frac{h_{i}-\mu}{\sqrt{\sigma^{2}+\epsilon}}$
@@
(9) $y_{i}=\gamma\hat{h}_{i}+\beta$
3.3. The FGAT Framework
@@
(10) $P_{r}^{link}(x,y)=\text{Sigmoid}(h_{x}h_{y}^{T})$
@@
Table 1. Experiment Results
Methods Ca-netscience Ca-sandi-auths
Precision Recall F1 ROC Precision Recall F1 ROC
MLP 0.4931 0.5879 0.5363 0.5044 0.5581 1.0000 0.7164 0.6181
GCN 0.5903 0.7363 0.6553 0.7078 0.5417 0.5417 0.5417 0.5304
GraphSAGE 0.5502 0.8132 0.6563 0.6277 0.4651 0.8333 0.5970 0.5885
GAT 0.6034 0.7692 0.6763 0.6916 0.6053 0.9583 0.7419 0.7240
FGAT 0.6667 0.6593 0.6630 0.7422 0.6216 0.9583 0.7541 0.7170
4. Experiments
@@
4.1. Datasets
Our evaluation utilizes two research collaboration network datasets, summarized in Table 2.
Table 2. Datasets Summary
@@ 4.1. Datasets
4.2. Experiment Settings and Evaluation Metrics
We benchmark FGAT against several prominent baseline models: MLP, GCN (kipf2016semi, ), GraphSAGE (hamilton2017inductive, ), and GAT (velivckovic2017graph, ). All baseline models maintain their default parameter configurations. For FGAT implementation, we configure the embedding dimension to 128 and employ a stack of 4 FGAT convolution layers. The dataset partitioning follows the aforementioned 0.7:0.1:0.2 ratio for training, validation, and testing, respectively.
@@ 4.2. Experiment Settings and Evaluation Metrics
4.3. Results
The experimental results are presented in Table 1, yielding several significant observations: